OpenStudy (anonymous):

I found the answer but need to make sure i've done the right steps to get there. Here is the problem i must answer: Sixty-Four can be split into four parts so that when 3 is added to the first part, 3 is subrtacted from the second part, the third part is multiplied by 3, and the fourth part is divided by 3, all four parts will be equal. What are the four parts?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (asnaseer):

start by giving a symbol to each of the four parts, lets call them: a, b, c and d

OpenStudy (asnaseer):

you know that:\[a+b+c+d=64\tag{1}\]

OpenStudy (asnaseer):

you are also told that:\[a+3=b-3=3c=d/3\tag{2}\]

OpenStudy (anonymous):

my teacher gave us a equation x+3+x-3+3x+x/3=64

OpenStudy (asnaseer):

that does not match the question you wrote out above

OpenStudy (asnaseer):

"Sixty-Four can be split into four parts" - I translate this to be:\[a+b+c+d=64\] "so that when 3 is added to the first part, 3 is subrtacted from the second part, the third part is multiplied by 3, and the fourth part is divided by 3, all four parts will be equal" - I translated this to:\[a+3=b-3=3c=d/3\]

OpenStudy (asnaseer):

you can then use equation (2) to express b, c and d in terms of "a". Then substitute those into equation (1) and solve to find "a". Then you can solve to find b, c and d.